OpenStudy (anonymous):

RADICAL EXPRESSION (8+√10)(8-√10) NO DECIMAL ANSWERS

OpenStudy (tkhunny):

Eyeball problem: \(8^{2} - (\sqrt{10})^{2} = 64 - 10\)

OpenStudy (anonymous):

???

OpenStudy (tkhunny):

You should have this in your bag of tricks: \(a^{2} - b^{2} = (a+b)(a-b)\). Have you seen it?

OpenStudy (anonymous):

thats not the answe

OpenStudy (tkhunny):

You didn't answer my question.

OpenStudy (anonymous):

im talking about the eyeball problem and yes ive seen that before

OpenStudy (tkhunny):

1) I didn't finish, completely, hoping you might see the rest. 2) Excellent. I am glad you have seen it. This solves the problem. Just see it backwards as it applies to your problem statement.